# PixFoundation: Are We Heading in the Right Direction with Pixel-level Vision Foundation Models?

### v1

@@ 1 Introduction
There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation (Zhou et al., 2022; Minaee et al., 2021; Kirillov et al., 2023; Ravi et al., 2024), pixel-level visual grounding and reasoning (Rasheed et al., 2024; Lai et al., 2024), depth estimation (Yang et al., 2024) and tracking (Wang et al., 2023). The majority of these have been transformed with the emergence of foundation models (Bommasani et al., 2021), specifically multi-modal large language models (MLLMs) (Liu et al., 2023/; Dai et al., 2023). Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance (Lai et al., 2024). Recent models tried to address this gap (Zhang et al., 2024b, a), yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
@@
Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks (Tong et al., 2024b, a). Such benchmarks focus on challenging visual patterns and tasks such as counting and relative positioning. Nonetheless, these benchmarks did not evaluate the recent pixel-level MLLMs. In this work we aim to provide challenging vision centric benchmarks that are dedicated to evaluate these models. Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA) among other capabilities?”. Our findings show that the majority of pixel-level MLLMs still fall short in such challenging setting. While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their capabilities.
@@
There have been recent works showing training-free open vocabulary segmentation emerging from vision language models (Wang et al., 2024; Luo et al., 2024; Hajimiri et al., 2025). Concurrent work has specifically explored emerging grounding in MLLMs (Cao et al., 2024). We are inspired by the previous work to establish a baseline on our proposed benchmarks. However, unlike their method we focus on the second research question of “When does grounding emerge in MLLMs that are not trained with full grounding supervision?”. Previous efforts studying emergent segmentation and/or grounding assumed correspondence with the specific language tokens (i.e., nouns or noun phrases) of the objects of interest. However, our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the object. We show it mostly emerges in the last 40% of the output, and can coincide with concepts about object parts, position/location, color/appearance or context of these objects. Fig. 1 summarizes our two major research questions.
@@
In summary, our contributions include: (i) Providing pixel-level segmentation annotations and referring expression of the object of interest in the question for two vision centric benchmarks (Tong et al., 2024a, b), which are PixMMVP and PixCV-Bench. (ii) Benchmarking recent efforts in pixel-level MLLMs where we show that they degrade VQA capabilities. More importantly, some of them lag in visual grounding with respect to simple techniques of extracting the segmentation from vanilla MLLMs, i.e., MLLMs that are not trained for pixel-level grounding. (iii) We provide a simple mechanism for extracting segmentation from vanilla MLLMs, with an understanding of when grounding emerges. Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest. We call it PixFoundation, since it mines the foundation model for pixel-level understanding.
@@ 2 Related Work
Pixel-level Vision Foundation Models. There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) (Kirillov et al., 2023; Ravi et al., 2024). Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision (Luo et al., 2024; Hajimiri et al., 2025; Wang et al., 2024). Yet, they relied on earlier foundation models that did not incorporate the power of large language models. Combining large language models with vision has been extensively researched with pioneering works such as Llava (Liu et al., 2023/, 2024) and instruct-BLIP (Dai et al., 2023). Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision (Lai et al., 2024; Rasheed et al., 2024; Zhang et al., 2024a, b, a, b). However, these methods were lagging in their chat performance. Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs (Tong et al., 2024b, a). Hence, it is still unclear if the pixel-level grounding supervision helped to improve their ability on these challenging tasks or not. In this work, we focus on the previous question to have a better understanding of their performance. Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding (Cao et al., 2024). We rely on this method as our baseline, but unlike previous works we provide an insight on when grounding emerges in such MLLMs.
Benchmarking Multi-modal Large Language Models There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU (Yue et al., 2024)) and their pixel-level counter parts (e.g., refCOCO/+/g (Yu et al., 2016; Kazemzadeh et al., 2014)). These benchmarks have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding. Nonetheless, there have been various works that discussed the shortcomings of MLLMs. In one of these works they discussed specifically CLIP (Radford et al., 2021), which is used in various MLLMs as the visual backbone. They proposed a benchmark, MMVP (Tong et al., 2024b), that is focused on the visual aspects within a VQA task designed on CLIP blind pairs. More recently, CV-Bench (Tong et al., 2024a) focused on two major tasks that are vision focused which are counting and relative positioning. Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output. Nonetheless, they still provide quite challenging scenarios that can act as a strong benchmark for the pixel-level MLLMs counterpart. In this work, we extend these two benchmarks to augment it with pixel-level annotations and referring expressions that correspond to the object of interest within their VQA task. These help to understand if failures occurring in these benchmarks stem from grounding or other reasons.
@@ 3.1 Benchmarks
PixMMVP benchmark: We build upon the recently released MMVP (Tong et al., 2024b) which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images. We augment the aforementioned dataset and manually annotate each question with the corresponding object of interest referring expression, e.g. an elderly person or the butterfly’s feet. There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded. Examples include questions inquiring on the view direction of the camera which are not tied to a specific entity. Our manual referring expression annotations are as fine-grained as possible. These expressions correspond to what needs to be grounded in the image to answer the question. Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator (Dutta et al., 2016).
@@
PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench (Tong et al., 2024a). We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K (Zhou et al., 2017) and COCO (Lin et al., 2014)), which can be used in our proposed benchmark. However, the publicly released CV-Bench does not identify the objects in question and their corresponding segmentation. As such we use GPT-4o to parse the questions and identify the objects of interest automatically, followed by manual inspection and correction. Specifically, we collect the classes in each image from the corresponding dataset and construct a list of class choices “1. CLS1, 2. CLS2, …”. Then we prompt GPT-4o with the following, “Provide number only as an answer. Identify the objects of interest in the following question: QUESTION ? 1. CLS1, 2. CLS2, … ”. This provides us with the categories per question that highlights the objects of interest. While seemingly these are categorical annotations not referring expressions, certain scenarios in CV-Bench are different. Specifically, in the relative positioning task all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categorical annotations.
@@
Afterwards, we use the selected categories from GPT-4o to retrieve the corresponding segmentation mask/s per image. Furthermore, we use a custom annotation tool to manually filter the objects in the question, e.g. selecting only the object mask annotated by the red box when referred to it and filtering out the other instances for that same class. Another example that needs manual filtration when the class in question is a broader category than what is inquired upon, e.g., “Pendant Lamp” which is under the category of “Lamp” in ADE20K. In such a case we need to filter out the masks of other types such as “Table Lamp”. Moreover, we identified missing annotations in rare occasions that required additional intervention. Examples include transparent objects such as wine glasses which were partially occluded, yet the ground-truth answer of the CV-Bench question included that object in the count. As such, we use VGG annotator for these missing annotations that were not part of the original segmentation masks. We provide the final PixCV-Bench with referring expressions and their segmentation annotations that can be used to evaluate the grounding ability in relation to the original VQA task. Appendix A provides visual examples from our benchmarks.
@@ 3.2 A Pixel-level MLLMs Study
Pixel-level MLLMs Shortcomings. We highlight the failures for the current state-of-the-art pixel-level MLLMs through three probing techniques. First, we highlight the degraded performance in VQA from most of the pixel-level MLLMs that are trained with pixel-level grounding supervision. We use for the first prompt “IMGQUESTION? OPTION1 OPTION2…”, as shown in Figure 2. Certain pixel-level MLLMs tend to answer the aforementioned question while outputting a corresponding segmentation mask/s for the objects of interest. Notably, the worst two models in this task, LISA (Lai et al., 2024) and GLAMM (Rasheed et al., 2024), are not able to provide an answer and rather refer to a segmentation mask. On the other hand, OMG-Llava (Zhang et al., 2024b) shows better ability in visual question answering.
@@
The second shortcoming we discuss is their degraded ability to visually ground objects. Surprisingly, although they were trained with pixel-level grounding supervision, not all of these models show superior grounding performance. Figure 3 shows the second prompt to generate a segmentation mask for the ground-truth referring expression. The purpose of this probing is to understand whether the failure in these models is purely on the VQA task, or its inability to ground the objects of interest in the corresponding question or both. Figure 3 shows the worst two models in this aspect, which are GLAMM, the region captioning variant, and Llava-G. Both fail to segment the specific object in question, while OMG-Llava shows better performance in grounding.
@@
Third, we highlight another shortcoming, where these MLLMs exhibit degraded ability to follow instructions. In order to probe this, we use the following prompt: “IMGQUESTION? a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA (Lai et al., 2024) and Llava-G (Zhang et al., 2024a). Both are incapable of following the instruction, yet Llava-G tries to tackle the question unlike LISA. On the other hand, OMG-Llava shows better ability to follow the instruction and answer the question.
@@
Baselines and upper bounds. In addition to evaluating state-of-the-art pixel-level MLLMs, we propose two baselines and one upper bound. The first of which is inspired by a concurrent work (Cao et al., 2024) that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision. Specifically, we use their attend and segment meta architecture as one of our baselines. However, we are the first to discuss when does such grounding emerge in these models. We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights on how these models reason.
@@
The attend and segment meta-architecture extracts the raw attention map for the output token, , where are the number of layers and heads, resp. Then, are the number of input language tokens before and after the visual tokens respectively, while are the height and width of the input image. Only the attention corresponding to the visual tokens of length are used, and these attention maps are averaged across the layers and heads, resulting in . This is further normalized across all the output, for output tokens. The attend and segment depends on using spaCy natural language processing tool (Honnibal et al., 2020) to identify the noun phrases and associate them to the ground-truth referring expressions. Thus, the spaCy embeddings closest to the ground-truth expression are used in the mask selection. This is followed by extracting the maximum attention point to feed into SAM (Kirillov et al., 2023) as a point prompt.
| Method | PixGr. Train | MMVP & PixMMVP | ||||
| Llava 1.5 (7B) (Liu et al., 2023/) | ✗ | 28.7 | 28.0 | - | - | - |
| Llava 1.5 (13B) (Liu et al., 2023/) | ✗ | 39.3 | 30.0 | - | - | - |
| Cambrian (8B)* (Tong et al., 2024a) | ✗ | 52.0 | 52.0 | - | - | - |
| OMG Llava (7B) (Zhang et al., 2024b) | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 |
| GLAMM (7B) (Rasheed et al., 2024) | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 |
| GLAMM - RegCap (7B) (Rasheed et al., 2024) | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 |
| LISA (7B) (Lai et al., 2024) | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 |
| Llava-G (7B) (Zhang et al., 2024a) | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 |
| Llava 1.5 (7B) + (a+s) (Cao et al., 2024) | ✗ | 28.7 | 28.0 | 11.1 | 11.2 | 16.1 |
| Llava 1.5 (13B) + (a+s) (Cao et al., 2024) | ✗ | 39.3 | 30.0 | 9.8 | 11.4 | 17.7 |
| Cambrian (8B)* + (a+s) (Cao et al., 2024) | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 |
| PixFoundation (Llava (7B)) (Ours) | ✗ | 28.7 | 28.0 | 16.9 | 18.8 | 22.7 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 39.3 | 30.0 | 14.4 | 18.2 | 24.9 |
| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 52.0 | 52.0 | 17.2 | 18.9 | 27.7 |
| Upper Bound - Oracle Selection | ||||||
| PixFoundation (Llava (7B)) (Ours) | ✗ | 28.7 | 28.0 | 26.1 | 38.0 | 32.7 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 39.3 | 30.0 | 23.6 | 38.2 | 38.7 |
| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 |
@@
For our baseline and upper bound, we build upon the previous pipeline and build an oracle upper bound and an automatic baseline. We introduce two main modifications to account for our observation that the correct grounding can occur with different output tokens describing the object not necessarily aligning with the exact ground-truth expression. The first modification is to go through all the potential output tokens without relying on spaCy embeddings. In the oracle we rely on the ground-truth mask to select the correct token and its corresponding attention map with highest intersection over union as an upper bound. The automatic baseline uses a simple but powerful mechanism where we overlay the predicted masks on the original image to highlight the potential object of interest. This is followed by feeding these images to a multi-modal LLM inquiring on which is best in highlighting this object. Specifically, we use the following prompt “Select the image that has EXPR best highlighted in red color than the others? Answer with a number from 1 to N and mention the number only. IMG”, where EXPR and IMG are the ground-truth expression and the image tokens respectively. In our experiments we rely on the open source Cambrian-1(8B) (Tong et al., 2024a) model in our automatic baseline for the mask selection. The second modification since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of a single one. This enables us to utilize the power of SAM in identifying fine-grained objects and referring expressions that tends to surpass what other MLLMs do, even those trained with pixel-level grounding supervision.
| Method | PixGr. Train | CV-Bench & PixCV-Bench | ||||
| Llava 1.5 (7B) (Liu et al., 2023/) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | - | - | - |
| Llava 1.5 (13B) (Liu et al., 2023/) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | - | - | - |
| Cambrian (8B)* (Tong et al., 2024a) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | - | - | - |
| OMG Llava (7B) (Zhang et al., 2024b) | ✓ | 9.2/14.7/12.0 | 36.8/47.4/42.1 | - | 50.5 | 45.9 |
| GLAMM (7B) (Rasheed et al., 2024) | ✓ | - | - | 30.2 | 51.9 | - |
| GLAMM - RCap (7B) (Rasheed et al., 2024) | ✓ | 22.8/32.8/27.8 | 46.8/62.0/54.4 | 3.6 | 7.4 | 13.0 |
| LISA (7B) (Lai et al., 2024) | ✓ | 1.9/5.5/3.7 | - | 16.8 | 48.1 | 6.7 |
| Llava-G (7B) (Zhang et al., 2024a) | ✓ | 13.9/14.2/14.1 | 5.1/3.7/4.4 | 1.7 | 17.6 | 15.8 |
| Llava 1.5 (7B) + (a+s) (Cao et al., 2024) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 5.2 | 15.6 | 24.8 |
| Llava 1.5 (13B) + (a+s) (Cao et al., 2024) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 4.7 | 14.9 | 24.0 |
| Cambrian (8B)* + (a+s) (Cao et al., 2024) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 18.6 | 15.9 | 29.6 |
| PixFoundation (Llava (7B)) (Ours) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 5.3 | 19.1 | 29.0 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 4.7 | 17.7 | 27.5 |
| PixFoundation (Cambrian (8B)*)(Ours) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 12.1 | 16.6 | 27.0 |
| Upper Bound - Oracle Selection | ||||||
| PixFoundation (Llava (7B)) (Ours) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 6.3 | 49.7 | 54.5 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 5.3 | 50.6 | 55.4 |
| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 54.3 | 64.4 | 68.1 |

### v2

@@ 1 Introduction
There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation [29, 14, 8, 17], pixel-level visual grounding and reasoning [16, 9], depth estimation [23] and tracking [21]. The majority of these have been transformed with the emergence of foundation models [1], specifically multi-modal large language models (MLLMs) [11, 3]. Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance [9]. Recent models tried to address this gap [27, 26], yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
@@
Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks [19, 18]. Such benchmarks focused on challenging visual patterns and tasks such as counting and relative positioning. Nonetheless, these benchmarks did not evaluate the recent pixel-level MLLMs. In this work we propose challenging vision centric benchmarks that are dedicated to evaluate these models. Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA)?”. Our findings show that the majority of pixel-level MLLMs still fall short in such a challenging setting. While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their chat capabilities.
@@
There have been recent works showing training-free open vocabulary segmentation emerging from vision language models [20, 13, 5]. Concurrent work has specifically explored emerging grounding in MLLMs [2]. We are inspired by the previous work to establish a baseline on our proposed benchmarks. However, unlike their method we focus on the second research question of “When does grounding emerge in MLLMs that are not trained with full grounding supervision?”. Previous efforts studying emergent segmentation and/or grounding assumed correspondence with the specific language tokens (i.e., nouns or noun phrases) of the objects of interest. However, our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the object. We show the most frequent emergence in PixMMVP happens in the last 40-60% of the output, and can coincide with concepts about object parts, position, color or context of these objects. Fig. 1 summarizes our research questions.
@@
In summary, our contributions include: (i) Proposing pixel-level vision centric benchmarks, PixMMVP and PixCV-Bench, with segmentation annotations and referring expression of the object of interest in the corresponding questions [18, 19]. (ii) Benchmarking recent efforts in pixel-level MLLMs where we show that they degrade VQA capabilities. More importantly, some of them lag in visual grounding with respect to simple techniques of extracting the segmentation from vanilla MLLMs, i.e., MLLMs that are not trained for pixel-level grounding. (iii) We provide a simple mechanism for extracting segmentation from vanilla MLLMs, with an understanding of when grounding emerges. Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest. We call it PixFoundation, since it can be plugged with any MLLM to mine for pixel-level understanding.
@@ 2 Related work
Pixel-level Vision Foundation Models. There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) [8, 17]. Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision [13, 5, 20]. Yet, they relied on earlier foundation models that did not incorporate the power of large language models. Combining large language models with vision has been extensively researched with pioneering works such as Llava [11, 12] and instruct-BLIP [3]. Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision [9, 16, 26, 27, 26, 27]. However, these methods were lagging in their chat performance. Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs [19, 18]. Hence, it is still unclear if the pixel-level grounding supervision helped to improve their ability on these challenging tasks or not. In this work, we focus on the previous question to have a better understanding of their performance. Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding [2]. We rely on this method as our baseline, but unlike previous works we provide an insight on when grounding emerges in such MLLMs. We propose another baseline that uses a novel and simple mechanism to perform mask selection while taking the previous insight into consideration. Our baseline outperforms the previous methods trained with full grounding supervision and the ones without it.
Benchmarking Multi-modal Large Language Models. There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU [25]) and pixel-level benchmarks (e.g., refCOCO/+/g [24, 7]). These have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding. Nonetheless, there have been various works that discussed the shortcomings of MLLMs. One of them discussed the shortcomings in CLIP [15], which is used in various MLLMs as a visual backbone. They proposed a benchmark, MMVP [19], that is focused on the visual aspects within a VQA task. More recently, CV-Bench [18] focused on two major tasks that are vision focused which are counting and relative positioning. Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output. Nonetheless, they still provide quite challenging scenarios that can act as a strong benchmark for the pixel-level MLLMs counterpart. In this work, we extend these two benchmarks to augment it with pixel-level annotations and referring expressions that correspond to the object of interest within their VQA task. These help to understand if failures occurring in these benchmarks stem from grounding or VQA capabilities.
@@ 3.1 Benchmarks
PixMMVP benchmark: We build upon the recently released MMVP [19] which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images. We augment the aforementioned dataset and manually annotate each question with the corresponding object of interest referring expression, e.g. an elderly person or the butterfly’s feet. There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded. Examples include questions inquiring on the view direction of the camera which are not tied to a specific entity. Our manual referring expression annotations are as fine-grained as possible. These expressions correspond to what needs to be grounded in the image to answer the question. Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator [4].
@@
PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench [18]. We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K [28] and COCO [10]), which can be used in our proposed benchmark. However, the publicly released CV-Bench does not identify the objects in question and their corresponding segmentation. As such we use GPT-4o to parse the questions and identify the objects of interest automatically, followed by manual inspection and correction. Specifically, we collect the classes in each image from the corresponding dataset and construct a list of class choices “1. CLS1, 2. CLS2, …”. Then we prompt GPT-4o with the following, “Provide number only as an answer. Identify the objects of interest in the following question: QUESTION ? 1. CLS1, 2. CLS2, … ”. This provides us with the categories per question that highlights the objects of interest. While seemingly these are categorical annotations not referring expressions, certain scenarios in CV-Bench are different. Specifically, in the relative positioning task all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categorical annotations.
@@
Afterwards, we use the selected categories from GPT-4o to retrieve the corresponding segmentation mask/s per image. Furthermore, we use a custom annotation tool to manually filter the objects in the question, e.g. selecting only the object mask annotated by the red box when referred to it and filtering out the other instances for that same class. Another example that needs manual filtration when the class in question is a broader category than what is inquired upon, e.g., “Pendant Lamp” which is under the category of “Lamp” in ADE20K. In such a case, we filter out the masks of other types such as “Table Lamp”. Moreover, we identify missing annotations in rare occasions that require additional intervention and manually annotate these missing objects. We provide the final PixCV-Bench with referring expressions and their segmentation annotations that can be used to evaluate the grounding ability in relation to the original VQA task. Appendix A provides visual examples from our benchmarks.
@@ 3.2 A Pixel-level MLLMs study
Pixel-level MLLMs shortcomings. We highlight the failures for the current state-of-the-art pixel-level MLLMs through three probing techniques. First, we highlight the degraded performance in VQA from most of the pixel-level MLLMs that are trained with pixel-level grounding supervision. We use for that the following prompt, “IMGQUESTION? OPTION1 OPTION2…”, as shown in Figure 2. Certain pixel-level MLLMs tend to answer the aforementioned question while outputting a corresponding segmentation mask/s for the objects of interest. Notably, the worst two models in this task, LISA [9] and GLAMM [16], are not able to provide an answer and rather refer to a segmentation mask. On the other hand, OMG-Llava [27] shows better ability in VQA.
@@
The second shortcoming we discuss is their degraded ability to visually ground objects. Surprisingly, although they were trained with pixel-level grounding supervision, not all of these models show superior grounding performance. Figure 3 shows the second prompt to generate a segmentation mask for the ground-truth referring expression. The purpose of this probing is to understand whether the failure in these models is purely on the VQA task, or its inability to ground the objects of interest in the corresponding question or both. Figure 3 shows the worst two models in this aspect, which are GLAMM, the region captioning variant, and Llava-G. Both fail to segment the specific object in question, while OMG-Llava shows better performance.
@@
Third, we highlight another shortcoming, where these MLLMs exhibit degraded ability to follow instructions. In order to probe this, we use the following prompt: “IMGQUESTION? a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA [9] and Llava-G [26]. Both are incapable of following the instruction, yet Llava-G tries to tackle the question unlike LISA. On the other hand, OMG-Llava shows better ability to follow the instruction and answer the question.
@@
Baselines and upper bounds. In addition to evaluating state-of-the-art pixel-level MLLMs, we propose two baselines and one upper bound. The first of which is inspired by a concurrent work [2] that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision. Specifically, we use their attend and segment meta architecture as one of our baselines. However, we are the first to discuss when does such grounding emerge in these models. We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights on how these models reason.
@@
The attend and segment meta-architecture extracts the raw attention map for the output token, , where are the number of layers and heads, resp. Then, are the number of input language tokens before and after the visual tokens respectively, while are the height and width of the input image. Only the attention corresponding to the visual tokens of length are used, and these attention maps are averaged across the layers and heads, resulting in . This is further normalized across all the output, for output tokens. The attend and segment depends on using spaCy natural language processing tool [6] to identify the noun phrases and associate them to the ground-truth referring expressions. Thus, the spaCy embeddings closest to the ground-truth expression are used in the mask selection. This is followed by extracting the maximum attention point to feed into SAM [8] as a point prompt.
@@
For our baseline and upper bound, we build upon the previous pipeline and build an oracle upper bound and an automatic baseline. We introduce two main modifications to account for our observation that the correct grounding can occur with different output tokens describing the object not necessarily aligning with the exact ground-truth expression. The first modification is to go through all the potential output tokens without relying on spaCy embeddings. In the oracle we rely on the ground-truth mask to select the correct token and its corresponding attention map with highest intersection over union as an upper bound. The automatic baseline uses a simple but powerful mechanism where we overlay the predicted masks on the original image to highlight the potential object of interest. This is followed by feeding these images to a multi-modal LLM inquiring on which is best in highlighting this object. Specifically, we use the following prompt “Select the image that has EXPR best highlighted in red color than the others? Answer with a number from 1 to N and mention the number only. IMG”, where EXPR and IMG are the ground-truth expression and the image tokens respectively. In our automatic baseline we rely on GPT-4o for the mask selection. The second modification, since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of one only. This enables us to utilize the power of SAM in identifying fine-grained objects and referring expressions that tends to surpass what other MLLMs do, even those trained with pixel-level grounding supervision.
| Method | PixGr. | PixMMVP | PixCV-Bench | ||||||||
| Llava 1.5 (7B) [11] | ✗ | 27.3 | 28.0 | - | - | - | 17.4 | 60.3 | - | - | - |
| Llava 1.5 (13B) [11] | ✗ | 39.3 | 30 | - | - | - | 14.5 | 61.4 | - | - | - |
| Cambrian (8B)* [18] | ✗ | 52.0 | 52.0 | - | - | - | 62.2 | 72.2 | - | - | - |
| OMG Llava (7B)** [27] | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 | 12.0 | 42.1 | - | 50.5 | 45.9 |
| GLAMM (7B) [16] | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 | - | - | 30.2 | 51.9 | - |
| GLAMM - RegCap (7B) [16] | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 | 27.8 | 54.4 | 3.6 | 7.4 | 13.0 |
| LISA (7B) [9] | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 | 3.7 | - | 16.8 | 48.1 | 6.7 |
| Llava-G (7B) [26] | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 | 14.1 | 4.4 | 1.7 | 17.6 | 15.8 |
| Llava 1.5 (7B) + (a+s) [2] | ✗ | 27.3 | 28.0 | 11.1 | 11.2 | 16.0 | 17.4 | 60.3 | 5.2 | 15.7 | 24.9 |
| Llava 1.5 (13B) + (a+s) [2] | ✗ | 39.3 | 30 | 9.8 | 11.4 | 17.7 | 14.5 | 61.4 | 4.7 | 14.9 | 24.0 |
| Cambrian (8B)* + (a+s) [2] | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 | 62.2 | 72.2 | 18.6 | 15.9 | 29.6 |
| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 18.8 | 25.9 | 26.9 | 17.4 | 60.3 | 5.4 | 28.5 | 38.7 |
| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 16.9 | 25.0 | 30.6 | 14.5 | 61.4 | 4.8 | 27.6 | 38.1 |
| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 29.6 | 30.3 | 38.3 | 62.2 | 72.2 | 23.9 | 33.1 | 45.4 |
| Upper Bound - Oracle Selection | |||||||||||
| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 26.1 | 38.0 | 32.2 | 17.4 | 60.3 | 6.3 | 49.7 | 54.5 |
| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 23.6 | 38.2 | 38.7 | 14.5 | 61.4 | 5.3 | 50.6 | 55.5 |
| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 | 62.2 | 72.2 | 54.3 | 64.4 | 68.1 |
@@
| Image | Prompt | Concept Category | Noun Phrase | Output |
[5pt]
1
[5pt]
3
[5pt]
6
[5pt]
161